# Exploring Learning Complexity for Efficient Downstream Dataset Pruning

### v2

@@ 5 Discussion
Does the proposed method depend on the quality of pre-trained models?
In the proposed masking-based learning complexity, the prediction error originates from the representation extracted from the pre-trained model, which generally shows efficacy in representation extraction (Radford et al., 2021). However, the representation quality can degrade when the pre-training dataset is defective, such as missing or noisy labels. To investigate the effectiveness of our method in this scenario, we utilize the ResNet-18 model pre-trained on the dataset with missing labels (Yalniz et al., 2019), to evaluate the downstream dataset pruning performance. As shown in Table 3, the proposed method outperforms the random baseline by 1.42% on average, keeping state-of-the-art performance. In general, DLC is not sensitive to the quality of pre-trained models. More experiment results with different architectures can be found in Appendix B.2.3.
Can the proposed scoring function transfer to pre-trained models with more parameters?
For better efficiency in scoring estimation, we transfer the proposed scoring function from a pre-trained ResNet-18 to models with more parameters. The pruning time cost and accuracy averaged over 5 downstream datasets with 9 pruning ratios are presented in Table 4. Despite consuming less time, we find that transferring does not achieve the original performance. Notably, its improvement over the random by 0.32% implies the effectiveness of our method.
Table 3: Average classification accuracy (%) over 5 diverse downstream datasets and pruning ratios with weakly pretrained ResNet-18. Bold numbers are the optimal results, and underline numbers are suboptimal results.
@@ 6 Conclusion
In this paper, we propose Distorted-based Learning Complexity (DLC), a novel and straightforward hardness score without relying on training or fine-tuning. We construct the learning path by distorting (masking) the initial pre-training weights with different ratios and calculate the loss integration via the Monte Carlo method for efficiency. Additionally, we design a flexible under-sampling strategy with randomness, named FlexRand, which can adapt to different data regimes while avoiding severe distribution shift. We conduct extensive experiments on the downstream image and instruction datasets pruning benchmarks, and the results show that our method establishes state-of-the-art performance. Notably, our method significantly speeds up pruning by 35$\times$ in the visual benchmark.
@@
Our method achieves the learning complexity by keeping multiple masked models with different masked rates. While the method does not require the computational cost of optimization, it may consume extra storage or memory to store these masked models.
Appendix A Related Works.
@@
To improve the robustness of different dataset pruning methods, the concept of Moderate Coreset (Xia et al., 2023) is discussed. Specifically, given any score criterion of data selection, different scenarios prefer data points with scores in different intervals.
However, the above-predefined scoring functions introduce additional pruning costs due to the parameter updating, which is not negligible (Qin et al., 2023). To reduce downstream dataset pruning costs, we propose an efficient scoring function named DLC solely based on the pre-trained model without fine-tuning.
Appendix B Experiments Details
B.1 Setups
B.1.1 Downstream Image Dataset Pruning Benchmark
@@
Aside from the random, we also compare the following methods: Herding (Welling, 2009), kCG (Sener & Savarese, 2018), CD (Agarwal et al., 2020), Forgetting (Toneva et al., 2018), Least Conf, Entropy, Margin (Coleman et al., 2020), GraNd, EL2N [^4] We perform three full fine-tuning with different seeds to estimate EL2N score accurately. (Paul et al., 2021), and SSP (Sorscher et al., 2022). If not specified, one-shot fine-tuning on the full downstream dataset is necessary to estimate the above scoring functions. For the masking-based learning complexity, the size of the learning path increases gradually, starting from the one with initial pre-training weights, until the ranking correlation with the previous no longer changes. The hyper-parameter $y$ in different data regimes is selected from the range {0.05, 0.10, 0.15, …, 0.95} by the validation accuracy.
@@
We prune the downstream datasets at 9 pruning ratios, ranging from 10% to 90%, for a thorough verification and comparison. For example, we keep 10% of each category in the original dataset when the pruning ratio is 10%. The efficacy and efficiency of fine-tuning dataset pruning is evaluated by measuring the following metrics: (1) downstream accuracy at different pruning ratios. (2) total time for scoring estimation, under-sampling, and associated hyper-parameter tuning.
@@
B.1.2 Downstream Instruction Dataset Pruning Benchmark
@@
B.2 Results
B.2.1 More results for DLC
Figure 7: Ranking correlation between the loss integration over the optimization and masking process. The first to fifth rows represent the results on CXRB10 (Nodule), DeepWeeds (Parthenium), DTD (Sprinkled), FGVCAircraft (DHC-6), and Sketch (Revolver) datasets respectively. (a,d,g,j,m): Loss trends with the number of parameters. (b,e,h,k,n): Loss trends with the optimization time. (c,f,I,l,o): High ranking correlation coefficient with $\rho=\{\mathbf{0.54},\mathbf{0.46},\mathbf{0.65},\mathbf{0.68},\mathbf{0.55}\}$.
B.2.2 Detailed Main Results
@@
B.2.3 More Results for Discussion

### v3

@@ 5 Discussion
Is the proposed method affected by the quality of pretraining datasets?
While our method provides outstanding performances on models pre-trained by fully supervised learning, it is underexplored whether the proposed method can work well on those models pre-trained on low-quality data, i.e., weakly supervised learning. To investigate the effectiveness of our method in this scenario, we utilize the ResNet-18 model pre-trained on the dataset with missing labels (Yalniz et al., 2019). We present the classification accuracy of those pre-trained models by linear probing in Appendix C.2.5. As shown in Table 3, the proposed method keeps state-of-the-art performance and outperforms the random baseline by 1.42% on average, not affected by the quality of the pre-training dataset. We provide detailed results in Appendix C.2.3.
Can we use small models to select samples for large models?
For better efficiency in scoring estimation, we transfer the proposed scoring function from a pre-trained ResNet-18 to models with more parameters. The pruning time cost and accuracy averaged over 5 downstream datasets with 9 pruning ratios are presented in Table 4. Notably, its improvement over the random by 0.32% implies the effectiveness of our method. However, we find that transferring does not achieve the original performance while consuming less time.
Table 3: Average accuracy (%) over 5 downstream datasets and varying pruning ratios with weakly pre-trained ResNet-18. Bold and underline numbers are the optimal and suboptimal results.
@@ 6 Conclusion
In this paper, we propose Distorted-based Learning Complexity (DLC), a novel and straightforward hardness score without relying on training or fine-tuning. We construct the learning path by distorting (masking) the initial pre-training weights with different ratios and calculate the loss integral via the Monte Carlo method for efficiency. Additionally, we design a flexible under-sampling strategy with randomness, named FlexRand, which can adapt to different data regimes while avoiding severe distribution shift. We conduct extensive experiments on the downstream image and instruction datasets pruning benchmarks, and the results show that our method establishes state-of-the-art performance. Notably, our method significantly speeds up pruning by 35$\times$ in the visual benchmark.
@@
The success of our method relies on the capability of pre-trained models. Thus, our method may fail to improve the downstream performance if the pre-trained model cannot provide meaningful representations.
Acknowledgments
This research is supported by the Shenzhen Fundamental Research Program (Grant No.JCYJ20230807091809020) and the SUSTech-NUS Joint Research Program. Bingyi Jing is supported in part by the National Natural Science Foundation of China under grant 12371290. We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology, and the Collaborative Innovation Center of Novel Software Technology and Industrialization at Nanjing University for our research.
Appendix A Related Works
@@
To improve the robustness of different dataset pruning methods, the concept of Moderate Coreset (Xia et al., 2023) is discussed. Specifically, given any score criterion of data selection, different scenarios prefer data points with scores in different intervals. However, the above-predefined scoring functions introduce additional pruning costs due to the parameter updating, which is not negligible (Qin et al., 2023). To reduce downstream dataset pruning costs, we propose an efficient scoring function named DLC solely based on the pre-trained model without fine-tuning.
Appendix B Method Details
In this section, we provide a concrete formulation of the pre-training weights masking operation. Given the pre-training weights $\mathbf{\it{W}}\in\mathbb{R}^{n*m}$ and masking ratio $r\in[0,1)$, the masking matrix $\mathbf{\it{M}}\in\{0,1\}^{n*m}$ is constructed by:
$\mathbf{\it{M}}_{i,j}=\begin{cases}0,&\mathrm{if}\ |\it{W}_{i,j}|<\tau_{r}\\ 1,&\mathrm{if}\ |\it{W}_{i,j}|\geq\tau_{r}\end{cases}$
where $\tau_{r}$ is the $({n*m*r})$-th element in $\{W_{1},...,W_{n*m}\}$ sorted by L1 norm in ascending order. Finally, the masked pre-training weights $\mathbf{\it{\hat{W}}}$ can be formulated as:
$\mathbf{\it{\hat{W}}}=\mathbf{\it{W}}\circ\mathbf{\it{M}}.$
Appendix C Experiments Details
C.1 Setups
C.1.1 Downstream Image Dataset Pruning Benchmark
@@
Aside from the random, we also compare the following methods: Herding (Welling, 2009), kCG (Sener & Savarese, 2018), CD (Agarwal et al., 2020), Forgetting (Toneva et al., 2018), Least Conf, Entropy, Margin (Coleman et al., 2020), GraNd, EL2N [^4] We perform three full fine-tuning with different seeds to estimate EL2N score accurately. (Paul et al., 2021), and SSP (Sorscher et al., 2022). If not specified, one-shot fine-tuning on the full downstream dataset is necessary to estimate the above scoring functions. For the masking-based learning complexity, the size of the learning path increases gradually, starting from the one with initial pre-training weights, until the ranking correlation with the previous no longer changes. The hyperparameter $y$ in different data regimes is selected from the range {0.05, 0.10, 0.15, …, 0.95} by the validation accuracy.
@@
We prune the downstream datasets at 9 pruning ratios, ranging from 10% to 90%, for a thorough verification and comparison. For example, we keep 10% of each category in the original dataset when the pruning ratio is 10%. The efficacy and efficiency of fine-tuning dataset pruning is evaluated by measuring the following metrics: (1) downstream accuracy at different pruning ratios. (2) total time for scoring estimation, under-sampling, and associated hyperparameter tuning.
@@
C.1.2 Downstream Instruction Dataset Pruning Benchmark
@@
C.2 Results
C.2.1 More results for DLC
Figure 7: Ranking correlation between the loss integral over the optimization and masking process. The first to fifth rows represent the results on CXRB10 (Nodule), DeepWeeds (Parthenium), DTD (Sprinkled), FGVCAircraft (DHC-6), and Sketch (Revolver) datasets respectively. (a,d,g,j,m): Loss trends with the number of parameters. (b,e,h,k,n): Loss trends with the optimization time. (c,f,I,l,o): High ranking correlation coefficient with $\rho=\{\mathbf{0.54},\mathbf{0.46},\mathbf{0.65},\mathbf{0.68},\mathbf{0.55}\}$.
C.2.2 Detailed Main Results
@@
C.2.3 More Results for Weakly Pre-trained Models
@@
C.2.4 Comparison of Different Under-sampling Strategies
Table 14: Average classification accuracy (%) over 9 pruning ratios with the fully pre-trained ResNet-18 from different under-sampling strategies. Bold numbers are the optimal results.
Under-sampling Strategy CXRB10 DeepWeeds DTD FGVCAircraft Sketch Average
Random 29.88 89.65 61.01 39.69 58.63 55.77
DQ 30.07 89.49 61.17 40.34 58.82 55.98
FlexRand 31.81 90.33 62.42 40.64 60.19 57.08
In this section, we further compare different under-sampling strategies to verify the superiority of FlexRand. Specifically, the strategy in Dataset Quantization (Zhou et al., 2023b) dividing datasets into different bins and randomly sampling from each bin, alleviates distribution shift and is similar to ours. Differently, FlexRand divides the full dataset into two bins with different sizes: easy and hard bins. In this way, samples in different bins have different probabilities of being selected, which enables adaptation in different data regimes. Moreover, we empirically compare the performance of FlexRand and Dataset Quantization (DQ) using the same DLC score. The experiments are conducted with the fully pre-trained ResNet-18 model and we keep the same fine-tuning setting as in the manuscript. As shown in Table 14, FlexRand outperforms the DQ, verifying its superiority.
C.2.5 Performance of Pre-trained Models by Linear Probing
Table 15: Accuracy (%) of pre-trained models by linear probing on the full downstream dataset.
Dataset Weakly Supervised Fully Supervised
RN18 RN50 ViT-Small ViT-Base RN18 RN50 ViT-Small ViT-Base
CXRB10 26.10 30.70 38.70 35.10 25.40 27.10 35.80 36.70
DeepWeeds 80.88 87.38 97.06 98.13 77.38 80.50 95.94 94.31
DTD 67.87 74.95 77.61 78.51 61.70 66.12 75.69 71.76
FGVCAircraft 35.55 46.38 69.10 71.92 27.96 30.06 64.21 61.12
Sketch 58.00 63.05 79.70 81.95 50.28 51.43 76.50 74.93
As shown in Table 15, we present the classification accuracy of those pre-trained models by linear probing by linear probing on the full downstream dataset.